OpenStudy (anonymous):

The coordinates of point A are (-4, p) and the coordinates of point B are ( 2, -3 ). The mid-point of the line segment AB, has coordinates ( q, 1). (a) Find the value of (i) q (ii) p (b) Calculate the distance AB.

OpenStudy (anonymous):

So are you aware of mid point formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. So mid point of AB is (q,1) Hence (-4+2/2, p-3/2) = (q,1) => (-1, p-3/2) = (q,1) so q = -1, and p = 5. (on solving) Hence A = (-4, 5) B = (2,-3) U can use distance formula to solve for AB. Do it yourself.(the answer should be 10 units.)

OpenStudy (anonymous):

so, if the midpoint formula says to add the x coordinates together and divide by 2, then add the y coordinates and divide by 2 in order to get the mid point. So we have (-4+2)=-2, then we divide by 2 to get -1, giving us the x value for the midpoint. For the Y value, we have (P-3) all divided by 2= the y value of 1 (given by the coordinate (q,1)). So, (P-3)/2=1. Multiply both sides by 2, and add 3 to find that p=5

OpenStudy (anonymous):

thank you both so much for the answers, that really clears things up!